HOW MANY LITERS OF A 20% SALT SOLUTION MUST BE ADDED TO 8 LITERS OF A 5% SALT SOLUTION TO YIELD AN ALLOY TAHT IS A 10% SALT SOLUTION?
Answer by mananth(16946)

HOW MANY LITERS OF A 20% SALT SOLUTION MUST BE ADDED TO 8 LITERS OF A 5% SALT SOLUTION TO YIELD AN ALLOY TAHT IS A 10% SALT SOLUTION?
...
20% x liters
5% 8 liters
Mix 10% (x+8) liters
...
0.2x+0.05*8 = 0.1(x+8)
0.2x+0.4=0.1x+0.8
0.1x=0.8-0.4
0.1x=0.4
x= 4 liters
...
m.ananth@hotmail.ca